OpenStudy (jaylenb):

A cannonball is fired toward a very tall vertical building 800 m away with an initial velocity of 100 m/s at 36.9° above the horizontal. The ball will

OpenStudy (campbell_st):

seems to be a lot of info missing

OpenStudy (jaylenb):

it is just asking what the ball will do

OpenStudy (anonymous):

My guess is that the question asks where the ball lands. In that case, use\[R = \dfrac{u^2\sin (2\theta)}{g} \]But again, we need to know the height of the building. If it is level with the horizontal, then we're fine.

OpenStudy (jaylenb):

here are the choices it gives A) hit the building in 2.0 s. B) hit the building in 4.0 s. C) hit the building in 8.0 s. D) hit the ground before hitting the building. E) hit the building in 10 s.

OpenStudy (anonymous):

First, ensure that the range of the projectile (horizontal distance covered) is above 800 metres.

OpenStudy (jaylenb):

it does not say in the question